OpenStudy (anonymous):

Given b = 4 and c = -5, evaluate b² - c².

OpenStudy (anonymous):

4² -(-5)² 4² =16 -(-5)² =-25 16-(-25)=41

OpenStudy (anonymous):

4^2-(-5^2) when you subtract a negative ou add a positive 4^2+5^2 square both numbers 16+25 add 41